OpenStudy (anonymous):

In a simple scenario with only two factors of production, suppose that capital’s share of income is .4 and labor’s share is .6 and that annual growth rates of capital and labor are 6 and 2 percent, respectively. Assume there is no productivity (technical) change. a. At what rate does output grow? b. How long will it take for output to double? Show. c. Now suppose productivity (technology) grows at a rate of 2 percent. Recalculate your answers to a. and b. above.

OpenStudy (anonymous):

the rate of output growth = (0.4x6) + (0.6x2)= 2.4+1.2= 3.6% b. 70/3.6 = 19 yrs c. 2 + 3.6%= 5.6% d. 70/5.6 = 12.5 yrs